OpenStudy (tw101):

Please help I have tried every single one of these equations and I cant seem to get it right. The width of a rectangle is 4 less than one-half its length. If the perimeter is 70 cm, identify the equation for the problem, then use it to find the length. Select two answers. 2(2l)+2(-4)=70 25 cm 2(l +1/2l - 4)=70 26 cm 2(l + 4l - 4/2)=70 21 cm 35 cm 2(2l)+2(4)=70

Nnesha (nnesha):

`he width of a rectangle is 4 less than one-half its length` translate this into math equation

Nnesha (nnesha):

the**

OpenStudy (tw101):

w= 1/2L?

Nnesha (nnesha):

hmmm there is a mistake the width of a rectangle is `4 less than one-half` its length

OpenStudy (tw101):

oh sorry w=1/2l-4

Nnesha (nnesha):

alright and what's the equation we should use to find perimeter of rectangle ?

OpenStudy (tw101):

2(l +1/2l - 4)=70

OpenStudy (tw101):

I tried it but I cant get the right answer

Nnesha (nnesha):

alright no the original equation is \[\huge\rm P=2L +2w\] that suppose to be our 2nd equation replace p with given perimeter which is 70

Nnesha (nnesha):

and then use substitution method

OpenStudy (tw101):

substitute for l and w?

Nnesha (nnesha):

\[\rm\color{ReD}{ w=\frac{1}{2}L-4}\] first equation^\[\rm2L+2\color{ReD}{w}=70\] 2nd equation^ so replace w with first equation

OpenStudy (tw101):

2L+2(1/2L-4)=70?

Nnesha (nnesha):

correct solve for l

Nnesha (nnesha):

L**

OpenStudy (tw101):

ok one second:)

OpenStudy (tw101):

I got like 37.5...

Nnesha (nnesha):

looks like you did the same mistake i was doing few sec ago

Nnesha (nnesha):

show me how did you get that ?

OpenStudy (tw101):

haha okay at least i wasnt too far off

OpenStudy (tw101):

umm so I distributed the two and then added like terms which resulted in 2l+3-8=70 which turned into 2l=75 and then i divided...

Nnesha (nnesha):

and how did you get 3 there ??

Nnesha (nnesha):

well there is a mistake with distribution \[2(\frac{1}{2}L-4)\] 2 times 1/2l = ?

OpenStudy (tw101):

1

Nnesha (nnesha):

not just one oh well let me retype it 2 times 1/2L

OpenStudy (tw101):

I meant 1l

Nnesha (nnesha):

\[\rm 2l+2(\frac{1}{2}L-4)=70\] \[\rm 2L+1L-8=70\] you should get 1L instead of 3 now combine like terms that's it:=)

OpenStudy (tw101):

oooh okay. thank you so much!

Nnesha (nnesha):

np :=)) good work!

OpenStudy (tw101):

thank you! and thanks for being patient with me haha. @Nnesha

Nnesha (nnesha):

my pleasure!